OpenStudy (anonymous):

Find the solution set for y = 3 – 2x, given the replacement set {(1, 1), (0, 3), (1, -2), (-1, 5)}.

OpenStudy (loser66):

juat plug the pair of number into the equation, which satisfy the equation, it 's a solution for that. for example : (0,3) that means x =0 , y =3 . plug them into y = 3 -2x you have 3 = 3 - 2 * 0 or 3 =3 . make sense. so ( 0,3) is a solution do the same with the rest.